OpenStudy (anonymous):

2. (7b2 – 5b + 3) – (b2 + 8b – 6)

OpenStudy (callisto):

\[(7b^2 – 5b + 3) – (b^2 + 8b – 6)\]\[= 7b^2 – 5b + 3 – b^2 - 8b –(-6)\]\[= 7b^2 – 5b + 3 – b^2 - 8b +6\]\[= 7b^2– b^2 – 5b- 8b + 3 +6\]\[=...?\]